•You have been hired to value a new 25 year callable, convertible bond. The bond has a 4.8 percent coupon, payable annually. The conversion price is $9, and the stock currently sells for $3.21. The stock price is expected to grow at 11 percent per year. The bond is callable at $120, but, based on prior experience, it won't be calledunlessthe conversion value is $130. The required return on this bond is 8 percent. Par value of the bond is $100. What value would you assign to this bond?
Answer #1

Hello

Straight Bond:

N = 25, PMT = $4.8, FV = $100, r = 8%

P 4.8 100 + (108戸= $65.84 (1:08)25]

Current value of straight bond = $65.84

Conversion ratio: CR = FV/CP = 100/9

Conversion value: CV = P*CR = 3.21*(100/9) = $35.67

Current conversion price = $35.67

Bond will be called when conversion price is $130

35.67*1.11^t=130

1.11^t=\frac{130}{35.67}

t=\frac{ln(\frac{130}{35.67})}{ln(1.11)}=12.392 years

PV=\frac{4.8}{0.08}*[1-\frac{1}{(1.08)^{12.392}}]+\frac{130}{(1.08)^{12.392}}=\$86.97

Current value of convertible bond = $86.97

I hope this clears your query.

Feel free to comment if you still have any questions or need additional information.

Do give a thumbs up if you find this helpful.